# A complex-projected Rayleigh quotient iteration for targeting interior eigenvalues

### v2

@@ 3.3 Generalised Eigenvalue Problems
for $\bm{y}^{(k+1)}$ at each step $k=0,1,\dotsc$, see [22, p. 358] for details. The procedure for the generalised RQI is then given by modifying Algorithm 2.1 by replacing the linear solve in Step 3 with (3.13) and instead normalising with respect to $\|\bm{x}\|_{M}\coloneqq\sqrt{\bm{x}^{*}M\bm{x}}$. Convergence can again be tested using the norm of the residual, which is now given by $\bm{r}^{(k)}_{M}\coloneqq(A-\mu^{(k)}M)\bm{x}^{(k)}$.
@@ 4 Numerical Experiments
In this section, we present numerical results for six different test problems structured in three examples: the first example discusses the case of real symmetric $3\times 3$ matrices; the second example comprises a comparison of classic RQI and PRQI for four different test matrices; and the last example comes from a practical model for photonic crystal fibres. All of the examples use the simplified PRQI outlined in Algorithm 3.2 with the choice $\gamma^{(k)}=\|\bm{r}^{(k)}\|$ for $k=0,1,\dotsc$. This is mainly motivated by the first example since in this case PRQI is invariant under scaling and shifting of the matrix under consideration (see Proposition 4.1 below), which allows for a direct comparison with classic RQI. We use a residual-based stopping criterion everywhere, i.e, the iteration is stopped when the norm of the residual falls below $\textsc{tol}=10^{-11}$ in Example 1, below $\textsc{tol}=10^{-15}$ in Example 2 and below $\textsc{tol}=10^{-8}$ in Example 3. The code that was used to generate the results is available online at https://github.com/nilsfriess/PRQI-Examples .
@@
Let $A\in\mathbb{R}^{3\times 3}$ be symmetric. In [20], the authors explain that for classic RQI it suffices to consider matrices $A=\mathrm{diag}(-1,s,1)$ for $s\in(-1,1)$ to characterise the convergence behaviour for the real symmetric $3\times 3$-case completely. The argument is based on the fact that classic RQI is scale and shift invariant [20, Proposition 3.1]. This also holds four our method as the following Proposition shows.
@@
Let $A\in\mathbb{R}^{n\times n}$ be real symmetric and let $B=\alpha A+\beta I$, $\alpha\neq 0$, $\beta\in\mathbb{R}$. Then, PRQI applied to $A$ and PRQI applied to $B$ produce the same iterates $\bm{x}^{(k)}$, $k=1,2,\dotsc,$ when started with the same vector $\bm{x}^{(0)}$.
@@
[…]
$\mu^{(0)}=\frac{{(\bm{x}^{(0)})}^{\ast}A\bm{x}^{(0)}}{{(\bm{x}^{(0)})}^{\ast}\bm{x}^{(0)}}=\tau^{2}s$
A straightforward calculation then shows that the (unnormalised) iterate after one step of classic RQI is given by
$\bm{x}^{(1)}={\Big{(}\frac{\varepsilon}{-1-\tau^{2}s}\,,\ \frac{\tau}{(1-\tau^{2})s}\,,\ \frac{\varepsilon}{1-\tau^{2}s}\Big{)}}^{T}.$
We observe the following: for any $s\in(0,1)$, the first entry of $\bm{x}^{(1)}$ is small because the denominator is bounded away from zero and $\varepsilon\approx 0$. Since $\tau\approx 1$, the denominator of the second entry is small and $\bm{x}^{(1)}$ has a large component in the target direction. The behaviour of the last entry depends on the value of $s$. For a large eigenvalue gap ($s\approx 0$), the last entry is small and $\bm{x}^{(1)}$ is a good approximation of the wanted eigenvalue. If, however, $s\approx 1$, i.e., the eigenvalue gap is small, the denominator of the last entry is small and, depending on how small $\varepsilon$ is, this entry might also blow up. In this case, $\bm{x}^{(1)}$ has a strong component in the direction of the (unwanted) eigenvector $\bm{v}_{3}$. For $s\in(-1,0]$ analogous observations can be made, with the roles of the first
In PRQI, adding the imaginary shift $\gamma=\gamma^{(0)}$ results in
$\bm{x}^{(1)}={\Big{(}\frac{\varepsilon}{-1-\tau^{2}s-i\gamma}\,,\ \frac{\tau}{(1-\tau^{2})s-i\gamma}\,,\ \frac{\varepsilon}{1-\tau^{2}s-i\gamma}\Big{)}}^{T}.$
In absolute value, the entries are given by
$\big{|}x^{(1)}_{1}\big{|}$ $=\frac{\varepsilon}{\sqrt{(1+\tau^{2}s)^{2}+\gamma^{2}}},\ \big{|}x^{(1)}_{2}\big{|}=\frac{\tau}{\sqrt{(1-\tau^{2})^{2}s^{2}+\gamma^{2}}},\ \big{|}x^{(1)}_{3}\big{|}=\frac{\varepsilon}{\sqrt{(1-\tau^{2}s)^{2}+\gamma^{2}}}.$
In this case, for $s\approx 1$, we have approximately $|x^{(1)}_{2}|\approx\tau/\gamma$ and $|x^{(1)}_{3}|\approx\varepsilon/\gamma$, implying that we do not get a strong component in the direction of the unwanted eigenvector.
@@
and our goal is to compute some of the discrete eigenvalues in the gap between those bands. We will exploit the fact that the qualitative behaviour of the desired eigenfunctions is very different from that of the undesired ones. As depticted in Figure 4, the target eigenfunctions decay exponentially away from the left boundary while the eigenfunctions corresponding to spurious eigenvalues decay exponentially away from the right boundary, see [18].
@@
Figure 5: Part of the spectrum of the original and the perturbed Sturm-Liouville eigenvalue problem (4.2). The target eigenvalue in the original spectrum (red cross) is very close to the unwanted eigenvalues in the band (hatched region) and the spurious eigenvalue (bordered circle). The perturbed eigenvalues (blue triangles) are raised into the complex plane and only the target eigenvalue (blue cross) stays near the real line.
@@
Figures 5(b) and 5(c) are the eigenvectors computed by PRQI and classic RQI, respectively using this initial vector. PRQI converged to the target eigenvalue $\lambda_{\text{PRQI}}=0.53874$ (which is in the gap between the bands $J_{1}$ and $J_{2}$ in (4.3)), while classic RQI converged to an eigenvalue far away ($\lambda_{\text{RQI}}=26.8648$). For classic RQI to converge to a value in the first spectral gap it was necessary to also set the initial shift appropriately.
@@
To target other eigenvalues in the first spectral gap, we can vary $R$ and/or $n_{\text{osc}}$. While PRQI succesfully avoided convergence to a spurious eigenvalue in all our experiements, it sometimes converged to an eigenvalue in one of the spectral bands. Since these eigenvalues correspond to the essential spectrum of the original problem, the computed eigenvectors are not localized anywhere, i.e., they oscillate over the whole interval $[0,X]$. This allows us to implement a modified PRQI that early on detects convergence to an unwanted eigenvalue. Let $S\in[0,X]$ and define
@@
Table 1 summarises the results using this approach for a few exemple values of $n_{\text{osc}}$ and $R$, the number of oscillations and cutoff values of the initial vectors that we used. The column “Index” contains the index of the eigenvalue computed by the respective method within the spectrum of the discretised equation (4.5) counted from the smallest eigenvalue.
@@
1.5 35 -0.227061 22 7 25.063959 174 8
2 35 0.349875 23 8 36.440082 209 6
2.5 35 0.538745 24 8 43.496076 228 6
3 55 0.349875 23 7 34.340555 203 7
3.5 55 0.538745 24 7 46.251764 235 4
4 55 0.581339 26 7 45.060462 232 7
Table 1: Results of applying PRQI and classic RQI to the Sturm–Liouville problem for different initial vectors parametrised by the number of oscillations $n_{\text{osc}}$ and the cutoff value $R$ PRQI converged to one of the target eigenvalues for many different combinations of $n_{\text{osc}}$ and $R$ while classic RQI always converged to an eigenvalue far away. Furthermore, PRQI successfully skips the spurious eigenvalue $\lambda_{\text{sp}}\approx 0.56$ with Index 25, and it required on average only one iteration more than classic RQI.
@@ 5 Conclusion
As we have remarked after the proof of Theorem 3.2, the convergence result that we proved is in a certain sense too weak. In particular, it can not explain the improved behaviour observed in the numerical examples. One direction for future research would therefore be to improve the convergence theorem; in the discussion following the current proof we already remarked that this likely requires a more accurate estimation of the eigenvalues of the matrix ${A}-i\gamma^{(k)}\bm{x}^{(k)}{(\bm{x}^{(k)})}^{\ast}I$.
Other possible avenues for future work would be to consider more examples of the Sturm–Liouville type as in Example 3. In [24] classic RQI and Inverse Iteration are combined to confine the eigenvalue iterates to a given interval. In the Sturm–Liouville example, the endpoints of the spectral gaps can be computed using Floquet theory [11], so possibly the idea in [24] can be extended to include PRQI: the combination of classic RQI and Inverse Iteration would ensure that the resulting eigenvalue lies within the wanted gap while PRQI avoids convergence to a spurious eigenvalue. Combining classic RQI and PRQI could also be fruitful in other cases as a means to improve convergence speed by switching from PRQI to classic RQI (i.e., setting $\gamma=0$) once the error residual is sufficiently small.

### v3

@@ 3.3 Generalised Eigenvalue Problems
for $\bm{y}^{(k+1)}$ at each step $k=0,1,\dotsc$, see [22, p. 358] for details. The procedure for the generalised RQI is then given by modifying Algorithm 2.1 by replacing the linear solve in Step 3 with (3.12) and instead normalising with respect to $\|\bm{x}\|_{M}\coloneqq\sqrt{\bm{x}^{\ast}M\bm{x}}$. Convergence can again be tested using the norm of the residual, which is now given by $\bm{r}^{(k)}_{M}\coloneqq(A-\mu^{(k)}M)\bm{x}^{(k)}$.
@@ 4 Numerical Experiments
In this section, we present numerical results for six different test problems structured in three examples: the first example discusses the case of real symmetric $3\times 3$ matrices; the second example comprises a comparison of classic RQI and PRQI for four different test matrices; and the last example comes from a practical model for photonic crystal fibres. All of the examples use the simplified PRQI outlined in Algorithm 3.2. The first two examples use $\gamma^{(k)}=\|\bm{r}^{(k)}\|$ for $k=0,1,\dotsc$. This is mainly motivated by the first example since in this case PRQI is invariant under scaling and shifting of the matrix under consideration (see Proposition 4.1 below), which allows for a direct comparison with classic RQI. The last example uses the squared residual norm for the imaginary shift, i.e., $\gamma^{(k)}=\|\bm{r}^{(k)}\|^{2}$. We use a residual-based stopping criterion everywhere, i.e., the iteration is stopped when the norm of the residual falls below $\textsc{tol}=10^{-11}$ in Example 1, below $\textsc{tol}=10^{-15}$ in Example 2 and below $\textsc{tol}=10^{-8}$ in Example 3. The code that was used to generate the results is available online at https://github.com/nilsfriess/PRQI-Examples .
@@
In [20], the authors explain that for classic RQI it suffices to consider matrices $A=\mathrm{diag}(-1,s,1)$ for $s\in(-1,1)$ to characterise the convergence behaviour for the real symmetric $3\times 3$-case completely. The argument is based on the fact that classic RQI is scale and shift invariant [20, Proposition 3.1]. This also holds for our method in the case where $\gamma^{(k)}$ is given by norm of the residual $\bm{r}^{(k)}$. Note that it does not hold for squared residual norm, since it is not invariant under linear transformations as in the proposition below.
@@
Let $A\in\mathbb{R}^{n\times n}$ be real symmetric and let $B=\alpha A+\beta I$, $\alpha\neq 0$, $\beta\in\mathbb{R}$. Then, for $\gamma^{(k)}=\|\bm{r}^{(k)}\|$, PRQI applied to $A$ and PRQI applied to $B$ produce the same iterates $\bm{x}^{(k)}$, $k=1,2,\dotsc,$ when started with the same vector $\bm{x}^{(0)}$.
@@
For the $[1,2,1]$ test matrix we repeated the experiment with a total of $10^{5}$ randomly sampled initial vectors and counted the number of successful runs (i.e., the number of initial vectors that led to convergence to the target eigenpair). Here, we used the variant of PRQI that uses the squared residual norm for the magnitude of the imaginary shift $\gamma^{(k)}$. In Table 1 we report the portion of successful runs for different ranges of initial error angles. We also report the fraction of runs where (3.4) is satisfied for the initial Rayleigh quotient shift $\mu^{(0)}$, as well as the mean of the magnitude of the initial imaginary shift for PRQI.
Initial error angle Portion of successful runs (3.4) satisfied Mean $\gamma^{(0)}$
Classic RQI PRQI
80 – 90 0 % 6.05 % 0 % 1.35
70 – 80 0 % 31.16 % 0 % 1.31
60 – 70 0.57 % 92.45 % 0 % 1.20
50 – 60 10.85 % 100 % 0 % 1.03
40 – 50 64.94 % 100 % 0.01 % 0.78
30 – 40 95.15 % 100 % 21.87 % 0.47
0 – 30 100 % 100 % 99.30 % 0.26
Table 1: Fraction of $10^{5}$ random initial vectors that led to convergence to the target eigenpair of the Laplace matrix in Example 2 for different ranges of initial error angles.
@@
and our goal is to compute some of the discrete eigenvalues in the gap between those bands. We will exploit the fact that the qualitative behaviour of the desired eigenfunctions is very different from that of the undesired ones. As depicted in Figure 4, the target eigenfunctions decay exponentially away from the left boundary while the eigenfunctions corresponding to spurious eigenvalues decay exponentially away from the right boundary, see [18].
@@
Figure 5: Part of the spectrum of the original and the perturbed Sturm–Liouville eigenvalue problem (4.2). The target eigenvalue in the original spectrum (red cross) is very close to the unwanted eigenvalues in the band (hatched region) and the spurious eigenvalue (bordered circle). The perturbed eigenvalues (blue triangles) are raised into the complex plane and only the target eigenvalue (blue cross) stays near the real line.
@@
Figures 6(b) and 6(c) are the eigenvectors computed by PRQI and classic RQI, respectively using this initial vector. PRQI converged to the target eigenvalue $\lambda_{\text{PRQI}}=0.53874$ (which is in the gap between the bands $J_{1}$ and $J_{2}$ in (4.3)), while classic RQI converged to an eigenvalue far away ($\lambda_{\text{RQI}}=26.8648$). For classic RQI to converge to a value in the first spectral gap it was necessary to also set the initial shift appropriately.
@@
To target other eigenvalues in the first spectral gap, we can vary $R$ and/or $n_{\text{osc}}$. While PRQI successfully avoided convergence to a spurious eigenvalue in all our experiments, it sometimes converged to an eigenvalue in one of the spectral bands. Since these eigenvalues correspond to the essential spectrum of the original problem, the computed eigenvectors are not localised anywhere, i.e., they oscillate over the whole interval $[0,X]$. This allows us to implement a modified PRQI that early on detects convergence to an unwanted eigenvalue. Let $S\in[0,X]$ and define
@@
Table 2 summarises the results using this approach for a few exemplar values of $n_{\text{osc}}$ and $R$, the number of oscillations and cutoff values of the initial vectors that we used. The column “Index” contains the index of the eigenvalue computed by the respective method within the spectrum of the discretised equation (4.5) counted from the smallest eigenvalue. In this case, we used the squared residual norm for the imaginary shift in PRQI, i.e., $\gamma^{(k)}=\|\bm{r}^{(k)}\|^{2}$ (which corresponds to $q=1$ in Theorem 3.2).
@@
1.5 35 -0.22706 22 7 25.06396 174 8
2 35 -0.22706 22 10 36.44008 209 6
2.5 35 -0.41034 10 8 43.49608 228 6
3 55 -0.22706 22 9 34.34056 203 7
3.5 55 0.34988 23 9 46.25176 235 4
4 55 0.34988 23 8 45.06046 232 7
4.5 55 0.53874 24 8 59.01389 $>$250 5
5 55 0.58134 26 8 68.37970 $>$250 5
Table 2: Results of applying PRQI (using $\gamma^{(k)}=\|\bm{r}^{(k)}\|^{2}$ in the imaginary shift) and classic RQI to the Sturm–Liouville problem for different initial vectors parametrised by the number of oscillations $n_{\text{osc}}$ and the cutoff value $R$ PRQI converged to one of the target eigenvalues for many different combinations of $n_{\text{osc}}$ and $R$ while classic RQI always converged to an eigenvalue far away. Furthermore, PRQI successfully skips the spurious eigenvalue $\lambda_{\text{sp}}\approx 0.56$ with Index 25, and it required on average only about two iteration more than classic RQI.
@@ 5 Conclusion
Possible avenues for future work would be to consider more examples of the Sturm–Liouville type as in Example 3. In [24] classic RQI and Inverse Iteration are combined to confine the eigenvalue iterates to a given interval. In the Sturm–Liouville example, the endpoints of the spectral gaps can be computed using Floquet theory [11], so possibly the idea in [24] can be extended to include PRQI: the combination of classic RQI and Inverse Iteration would ensure that the resulting eigenvalue lies within the wanted gap while PRQI avoids convergence to a spurious eigenvalue. Combining classic RQI and PRQI could also be fruitful in other cases as a means to improve convergence speed by switching from PRQI to classic RQI (i.e., setting $\gamma=0$) once the error residual is sufficiently small.